# reference.bib

### v1

@@ 2. Preliminaries
We define the Heisenberg group to be:
@@
The group $SL_{2}(\mathbb{R})$ acts on $H^{(\Lambda,\sigma)}(\mathbb{R})$ from the right, via:
@@
The real Jacobi group associated with $(\Lambda,\sigma)$, denoted with $J^{(\Lambda,\sigma)}(\mathbb{R})$, is defined to be the semi-direct product of $SL_{2}(\mathbb{R})$ and $H^{(\Lambda,\sigma)}(\mathbb{R})$. The composition law is then:
@@
We also define the integral Jacobi group to be the semi-direct product of $SL_{2}(\mathbb{Z})$ and $H^{(\Lambda,\sigma)}(\mathbb{Z})$ and we will denote it by $J^{(\Lambda,\sigma)}$.
We are now going to define some slash operators, acting on holomorphic, complex-valued functions on ×$\mathbb{H}\times(\Lambda\otimes\mathbb{C})$.
@@
Let $k$ be a positive integer and $f:\mathbb{H}\times(\Lambda\otimes\mathbb{C})\longrightarrow\mathbb{C}$ a holomorphic function. For $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL_{2}(\mathbb{R})$ , we define:
$f|_{k,(\Lambda,\sigma)}[M](\tau,z):=(c\tau+d)^{-k}e^{-\pi ic\sigma(z,z)/(c\tau+d)}f\left(\frac{a\tau+b}{c\tau+d},\frac{z}{c\tau+d}\right).$
In the case when $M\in GL_{2}^{+}(\mathbb{R})$, we use $\det(M)^{-1/2}M$ instead of $M$. For $h=(x,y,\zeta)\in H^{(\Lambda,\sigma)}(\mathbb{R})$:
$f\mid_{k,(\Lambda,\sigma)}h(\tau,z):=\zeta\cdot e^{\pi i\tau\sigma(x,x)+2\pi i\sigma(x,z)}f(\tau,z+x\tau+y).$
Finally, for the action of $J^{(\Lambda,\sigma)}(\mathbb{R})$ on complex-valued, holomorphic functions on ×$\mathbb{H}\times(L\otimes\mathbb{C})$, we have:
$(f,(A,h))\longmapsto f\mid_{k,(\Lambda,\sigma)}(A,h)(\tau,z):=(f\mid_{k,(\Lambda,\sigma)}A)\mid_{k,(\Lambda,\sigma)}h.$
We now have the following definition ([jacobi_lattice, Definition 1.23]):
@@
Let $V_{\mathbb{C}}:=V\otimes\mathbb{C}$ and extend $\sigma$ to $V_{\mathbb{C}}$ by $\mathbb{C}$-linearity. For $k$ a positive integer, a holomorphic function $f:\mathbb{H}\times V_{\mathbb{C}}\rightarrow\mathbb{C}$ (where $\mathbb{H}$ denotes the usual upper half plane) is called a Jacobi form of weight $k$ with respect to $(\Lambda,\sigma)$ if the following hold:
  • For all $\gamma\in J^{(\Lambda,\sigma)}$, we have
    $f\mid_{k,(\Lambda,\sigma)}\gamma(\tau,z)=f(\tau,z).$
@@
    $f(\tau,z)=\sum_{n\in\mathbb{Z},r\in\Lambda^{*},2n\geq\sigma[r]}c_{f}(n,r)e(n\tau+\sigma(r,z)).$
[…]
We now have a notion of a scalar product for elements of $S_{k}(\Lambda,\sigma)$ ([jacobi_lattice, Definition 1.33]).
@@
Let $\phi,\psi\in S_{k}(\Lambda,\sigma)$. If then $U\leq J^{(\Lambda,\sigma)}$ of finite index, we define their Petersson inner product via:
[…]
where $\tau=u+iv$, $z=x+iy$ and $dV:=v^{-n-2}dudvdxdy$. This inner product does not depend on the choice of $U$, so in what follows, we drop the subscript.
From now on, we specify to our case, by taking $\Lambda=L=\mathbb{Z}^{n}$ and $\sigma(x,y)=x^{t}Sy$ for all $x,y\in V$.
[…]
$F(\tau^{\prime}+m,z,\tau)=F(\tau^{\prime},z,\tau)$
(by the form of elements belonging to $\Gamma_{S}$). Hence, we can write
$F(Z)=\sum_{m\geq 0}\phi_{m}(\tau,z)e^{2\pi im\tau^{\prime}},$
[…]
$\phi_{m}(\tau,z)=\sum_{l=0}^{\infty}\sum_{\begin{subarray}{c}\lambda\in\mathbb{Z}^{n},\\ 2ml\geq\lambda^{t}S^{-1}\lambda\end{subarray}}c_{\phi_{m}}(\lambda,l)e^{2\pi i(l\tau+\lambda^{t}z)}.$
We note that then $\phi_{m}\in J_{k}(\mathbb{Z}^{n},m\sigma)$.
We will now re-write the above Fourier expansion of a cusp form in a form similar to Kohnen-Skoruppa in [kohnen_skoruppa]. By the above, we have:
$\phi_{m}(\tau,z)=\sum_{n\in\mathbb{Z},r\in\mathbb{Z}^{n},2mn>r^{t}S^{-1}r}c_{\phi}(n,r)e(n\tau+r^{t}z).$
Now, from the condition $2mn>r^{t}S^{-1}r$ and the definition of the level, we can write
$mnq-\frac{1}{2}qS^{-1}[r]=-D\implies n=\frac{\frac{1}{2}qS^{-1}[r]-D}{qm},$
with some integer $D<0$.
So, in total, we get the following lemma:
Lemma 3.1.
[…]
$\phi_{m}(\tau,z)=\sum_{\begin{subarray}{c}D\in\mathbb{Z}_{<0},r\in\mathbb{Z}^{n}\\ \frac{1}{2}qS^{-1}[r]\equiv D\pmod{qm}\end{subarray}}c_{\phi_{m}}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qm}\tau+r^{t}z\right).$
@@
By adjusting [jacobi_lattice, Proposition 1.25] to the notation of the above Lemma 3.1, we have that
$c_{\phi_{m}}(D,r)=c_{\phi_{m}}(D^{\prime},r^{\prime})\iff D=D^{\prime}\textup{and }S^{-1}r\equiv S^{-1}r^{\prime}\pmod{m\mathbb{Z}^{n}}.$
[…]
Theorem 3.1.
With the notation as above, the mapping $F\longmapsto\phi_{1}$ gives an isomorphism between the Maass space $M_{k}^{*}(\Gamma_{S})$ and $J_{k}(L,\sigma)$, as vector spaces.
@@ 4. Poincaré Series
[…]
@@
We define the support of the lattice $L$ with respect to the bilinear form $\sigma(x,y)=x^{t}Sy$ for $x,y\in V$ to be
@@
which then implies $D\in\frac{1}{q}\mathbb{Z}$. Hence, by writing $D\longmapsto\frac{D}{q}$, we get that equivalently
@@
$\widetilde{\textup{supp}}(L,\sigma):=\{(D,r)\mid(D/q,S^{-1}r)\in\textup{supp}(L,\sigma)\}.$
@@
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. We then define the following complex valued function on ×$\mathbb{H}\times(L\otimes\mathbb{C})$:
$g_{D,r}(\tau,z):=e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}+r^{t}z\right).$
@@
We now have the following very important property, which can be found in [jacobi_lattice, Theorem 2.3, (i)].
@@
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. Then for any $f\in J_{k}(L,\sigma)$ with
@@
5. Operator $V_{N}$ and Dirichlet Series
In this section, we again denote by $\sigma(x,y)=x^{t}Sy$ for $x,y\in V$. Given a Jacobi form $\phi\in J_{k}(\mathbb{Z}^{n},\sigma)$ and $N\geq 1$, we define the operator:
$V_{N}:J_{k}(\mathbb{Z}^{n},\sigma)\longrightarrow J_{k}(\mathbb{Z}^{n},N\sigma),$
given by (compare also with [sugano_lifting, Section 6], [jacobi_lattice, Definition 4.25]):
$(V_{N}\phi)(\tau,z)=N^{k-1}\sum_{M\in SL_{2}(\mathbb{Z})\backslash M_{2}(\mathbb{Z})_{N}}(c\tau+d)^{-k}e^{-\pi icNS[z]/(c\tau+d)}\phi\left(M\langle\tau\rangle,\frac{Nz}{c\tau+d}\right),$
where $M_{2}(\mathbb{Z})_{N}=\{g\in M_{2}(\mathbb{Z})|\det g=N\}$ and for $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix},\tau\in\mathbb{H}$, we define
$M\langle\tau\rangle:=\frac{a\tau+b}{c\tau+d}.$
[…]
Proposition 5.1.
[…]
$\longmapsto\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod q\end{subarray}}\left(\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}s\mod dS\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[s]\mod{qd}\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}s\right)\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right).$
Remark 2.
The $(D,r)$ coefficient of $V_{N}^{*}\phi$ is independent of $r$.

### v2

@@ 3. Fourier–Jacobi Forms
We define the Heisenberg group to be
@@
The group $\hbox{SL}_{2}(\mathbb{R})$ acts on $H^{(\Lambda,\sigma)}(\mathbb{R})$ from the right, via
@@
The real Jacobi group associated with $(\Lambda,\sigma)$, denoted with $J^{(\Lambda,\sigma)}(\mathbb{R})$, is defined to be the semi-direct product of $\hbox{SL}_{2}(\mathbb{R})$ and $H^{(\Lambda,\sigma)}(\mathbb{R})$. The composition law is then
@@
We also define the integral Jacobi group to be the semi-direct product of $\hbox{SL}_{2}(\mathbb{Z})$ and $H^{(\Lambda,\sigma)}(\mathbb{Z})$ and we will denote it by $J^{(\Lambda,\sigma)}$.
We are now going to define some slash operators, acting on holomorphic, complex-valued functions on $\mathbb{H}\times(\Lambda\otimes\mathbb{C})$.
@@
Let $k$ be a positive integer and $f:\mathbb{H}\times(\Lambda\otimes\mathbb{C})\longrightarrow\mathbb{C}$ a holomorphic function. For $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\hbox{SL}_{2}(\mathbb{R})$ , we define
$(f|_{k,(\Lambda,\sigma)}[M])(\tau,z):=(c\tau+d)^{-k}e^{-\pi ic\sigma(z,z)/(c\tau+d)}f\left(\frac{a\tau+b}{c\tau+d},\frac{z}{c\tau+d}\right).$
In the case when $M\in\hbox{GL}_{2}^{+}(\mathbb{R})$, we define the action by using $\det(M)^{-1/2}M$ instead of $M$. For $h=(x,y,\zeta)\in H^{(\Lambda,\sigma)}(\mathbb{R})$, we define
$(f\mid_{k,(\Lambda,\sigma)}h)(\tau,z):=\zeta\cdot e^{\pi i\tau\sigma(x,x)+2\pi i\sigma(x,z)}f(\tau,z+x\tau+y).$
Finally, for the action of $J^{(\Lambda,\sigma)}(\mathbb{R})$, we have
$(f,(A,h))\longmapsto(f\mid_{k,(\Lambda,\sigma)}(A,h))(\tau,z):=(f\mid_{k,(\Lambda,\sigma)}A)\mid_{k,(\Lambda,\sigma)}h.$
We now have the following definition ([17, Definition 1.23]).
@@
Let $V_{\mathbb{C}}:=V\otimes\mathbb{C}$ and extend $\sigma$ to $V_{\mathbb{C}}$ by $\mathbb{C}$-linearity. For $k$ a positive integer, a holomorphic function $f:\mathbb{H}\times V_{\mathbb{C}}\rightarrow\mathbb{C}$ (where $\mathbb{H}$ denotes the usual upper half plane) is called a Fourier–Jacobi form of weight $k$ with respect to $(\Lambda,\sigma)$ if the following hold:
  • For all $\gamma\in J^{(\Lambda,\sigma)}$ and $(\tau,z)\in\mathbb{H}\times V_{\mathbb{C}}$, we have
    $(f\mid_{k,(\Lambda,\sigma)}\gamma)(\tau,z)=f(\tau,z).$
@@
    $f(\tau,z)=\sum_{m\in\mathbb{Z},r\in\Lambda^{*},2m\geq\sigma[r]}c_{f}(m,r)e(m\tau+\sigma(r,z)).$
We denote the space of such forms by $J_{k}(\Lambda,\sigma)$. We say $f$ is a Jacobi cusp form if $c_{f}(m,r)=0$ when $2m=\sigma[r]$. We denote the space of Fourier–Jacobi cusp forms by $S_{k}(\Lambda,\sigma)$.
Remark 1.
We could also characterise Fourier–Jacobi forms as modular forms transforming under the action of the parabolic subgroup of $\Gamma_{S}$, given by
$\Gamma_{S,J}:=\left\{\begin{pmatrix}*&*\\ 0&D\end{pmatrix}\in\Gamma_{S}\mid D\in\textup{SL}_{2}(\mathbb{Z})\right\}.$
This is true because there is an embedding $\iota:J_{S}\longrightarrow\Gamma_{S,J}$ (see [21, p. 44] or [28, p. 27]).
We now have a notion of a scalar product for elements of $S_{k}(\Lambda,\sigma)$ ([17, Definition 1.33]).
@@
Let $\phi,\psi\in S_{k}(\Lambda,\sigma)$. If then $U\leq J^{(\Lambda,\sigma)}$ of finite index, we define their Petersson inner product via
$\langle\phi,\psi\rangle_{U}=\frac{1}{\left[J^{\Lambda}:U\right]}\int_{U\backslash\mathcal{H}\times(\Lambda\otimes\mathbb{C})}\phi(\tau,z)\overline{\psi(\tau,z)}v^{k}e^{-2\pi\sigma(y,y)v^{-1}}\textup{d}V,$
where $\tau=u+iv$, $z=x+iy$ and $\textup{d}V:=v^{-n-2}\textup{d}u\textup{d}v\textup{d}x\textup{d}y$. This inner product does not depend on the choice of $U$, so in what follows, we drop the subscript.
We now specify to our case, by taking $\Lambda=L=\mathbb{Z}^{n}$ and $\sigma(x,y)=x^{t}Sy$ for all $x,y\in V$.
Let us discuss the Fourier–Jacobi expansion of orthogonal cusp forms of weight $k$ with respect to a subgroup $\widetilde{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$. If we write $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$ with $\tau^{\prime},\tau\in\mathbb{C},\ z\in\mathbb{C}^{n}$, we have that for $F\in S_{k}(\Gamma)$ and any $m\in\mathbb{Z}$:
$F(\tau^{\prime}+m,z,\tau)=F(\tau^{\prime},z,\tau).$
Hence, we can write
$F(Z)=\sum_{m\geq 1}\phi_{m}(\tau,z)e^{2\pi im\tau^{\prime}},$ (3)
and we call the functions $\phi_{m}(\tau,z)$ the Fourier–Jacobi coefficients of $F$. We note that then $\phi_{m}\in S_{k}(\mathbb{Z}^{n},m\sigma)$ (see [28, (5.10), (5.11)]).
4. $V_{N}$ operator
Let now again $\sigma(x,y)=x^{t}Sy$ for $x,y\in V$. Let $N\geq 1$ and define $M_{2}(\mathbb{Z})_{N}:=\{g\in M_{2}(\mathbb{Z})|\det g=N\}$. For any $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in M_{2}(\mathbb{Z})_{N}$ and $\tau\in\mathbb{H}$ (usual upper half plane), we define
$M\langle\tau\rangle:=\frac{a\tau+b}{c\tau+d}.$
Given a Fourier–Jacobi cusp form $\phi\in S_{k}(\mathbb{Z}^{n},\sigma)$, we define the operator
$V_{N}:S_{k}(\mathbb{Z}^{n},\sigma)\longrightarrow S_{k}(\mathbb{Z}^{n},N\sigma),$
given by
$(V_{N}\phi)(\tau,z)=N^{k-1}\sum_{M\in\textup{SL}_{2}(\mathbb{Z})\backslash M_{2}(\mathbb{Z})_{N}}(c\tau+d)^{-k}e^{-\pi icNS[z]/(c\tau+d)}\phi\left(M\langle\tau\rangle,\frac{Nz}{c\tau+d}\right),$
This is well-defined by [28, Lemma 6.1] or [17, Definition 4.25]. Our aim in this Section is to compute its adjoint with respect to the scalar product of Fourier–Jacobi forms. This is the analogue of the main Proposition in [13].
Now, if $\phi\in S_{k}(\mathbb{Z}^{n},N\sigma)$, we will write its Fourier expansion in a form similar to Kohnen and Skoruppa in [13, Section 2]. From Definition 3.4, we can write
$\phi(\tau,z)=\sum_{\begin{subarray}{c}m\in\mathbb{Z},r\in\mathbb{Z}^{n}\\ 2Nm>r^{t}S^{-1}r\end{subarray}}c_{\phi}(m,r)e(m\tau+r^{t}z),$
for some $c_{\phi}(m,r)\in\mathbb{C}$. Now, from the condition $2Nm>r^{t}S^{-1}r$ and the definition of the level (see Definition 2.4), we can write
$Nmq-\frac{1}{2}qS^{-1}[r]=-D\implies m=\frac{\frac{1}{2}qS^{-1}[r]-D}{qN},$
for some integer $D<0$. Therefore, we can write
$\phi(\tau,z)=\sum_{\begin{subarray}{c}D\in\mathbb{Z}_{<0},r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod qN\end{subarray}}c_{\phi}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\tau+r^{t}z\right).$ (4)
@@ 4. $V_{N}$ operator
By adjusting [17, Proposition 1.25] to the notation above, we have that
$D=D^{\prime}\textup{ and }S^{-1}r\equiv S^{-1}r^{\prime}\pmod{N\mathbb{Z}^{n}}\implies c_{\phi}(D,r)=c_{\phi}(D^{\prime},r^{\prime}).$
We are now ready to give the result concerning the adjoint of $V_{N}$. We use the notation that $s\equiv s^{\prime}\ (\mathrm{mod}\ NS\mathbb{Z}^{n})\iff s-s^{\prime}=NSu$ for some $u\in\mathbb{Z}^{n}$.
Proposition 4.1.
The action of $V_{N}^{*}$, defined as the adjoint of $V_{N}$ with respect to the scalar product of Fourier–Jacobi forms, defined in 3.5, is given by (for $\phi\in S_{k}(\mathbb{Z}^{n},N\sigma)$)
$\longmapsto\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod q\end{subarray}}\left(\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}s\mod dS\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[s]\mod{qd}\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}s\right)\right)\times\\ \times e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right).$
Remark 3.
The $(D,r)$ coefficient of $V_{N}^{*}\phi$ is independent of $r$.
Proof.
We start by writing
$V_{N}\phi=N^{k/2-1}\sum_{A\in\textup{SL}_{2}(\mathbb{Z})\backslash M_{2}(\mathbb{Z})_{N}}\phi_{\sqrt{N}}|_{k,(\mathbb{Z}^{n},N\sigma)}\left(\frac{1}{\sqrt{N}}A\right),$
where we define $\phi_{\sqrt{N}}(\tau,z):=\phi\left(\tau,\sqrt{N}z\right)$. We remind here that if $A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\textup{GL}_{2}^{+}(\mathbb{R})$, we use $(\det A)^{-1/2}A$ in the $\mid_{k}$-action defined in Definition 3.3.
We can then follow the proof in [13, pages 554-556] line by line and arrive at the result that the adjoint $V_{N}^{*}:S_{k}(\mathbb{Z}^{n},N\sigma)\longrightarrow S_{k}(\mathbb{Z}^{n},\sigma)$ is given by
$\phi\longmapsto N^{k/2-2n-1}\sum_{X\ (\mathrm{mod}\ N\mathbb{Z}^{2n})}\sum_{A\in\textup{SL}_{2}(\mathbb{Z})\backslash M_{2}(\mathbb{Z})_{N}}\phi_{\sqrt{N}^{-1}}|_{k,(\mathbb{Z}^{n},\sigma)}\left(\frac{1}{\sqrt{N}}A\right)|_{k,(\mathbb{Z}^{n},\sigma)}X.$
Let us now compute the action on the Fourier coefficients. We choose a set of representatives for $\textup{SL}_{2}(\mathbb{Z})\backslash M_{2}(\mathbb{Z})_{N}$ of the form $\begin{pmatrix}a&b\\ 0&d\end{pmatrix}$ with $ad=N$ and $0\leq b<d$. We then obtain that the above expression can be written as
By then substituting the Fourier expansion
for $\phi$, the above can be written as
Now, the term
is zero unless
$d\mid\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\textup{ and }d\mid r,$
meaning it divides all of its components. In that case, the sum equals $dN^{n}$. The conditions then imply that we can replace $r\longmapsto dr$ and $D\longmapsto Dd^{2}$. The last one follows from the fact that $N=ad$ and so $d\mid N$ as well. Hence, we obtain that the expression equals
$N^{k-n-1}\sum_{\lambda\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{d\mid N}d^{1-k}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\ (\mathrm{mod}\ \frac{qN}{d})\end{subarray}}c_{\phi}(d^{2}D,dr)\times\\ \times e\left(\left(\frac{\frac{1}{2}qS^{-1}[r]+\frac{1}{2}qS[\lambda]+qr^{t}\lambda-D}{q}\right)\tau+r^{t}z+\lambda^{t}Sz\right)=$
$N^{k-n-1}\sum_{d\mid N}d^{1-k}\sum_{\lambda\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r-S\lambda]\ (\mathrm{mod}\ \frac{qN}{d})\end{subarray}}c_{\phi}(d^{2}D,d(r-S\lambda))\times\\ \times e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right),$ (5)
after setting $r\longmapsto r-S\lambda$. Now, as in [13, p. 557], we set
$\lambda\equiv t+\frac{N}{d}t^{\prime}\ (\mathrm{mod}\ N\mathbb{Z}^{n}),$ (6)
with $t\ (\mathrm{mod}\ (N/d)\mathbb{Z}^{n})$ and $t^{\prime}\ (\mathrm{mod}\ d\mathbb{Z}^{n})$. We then have
$S^{-1}(d(r-S\lambda))\equiv S^{-1}(d(r-St))\mod N\mathbb{Z}^{n},$
and
$D\equiv\frac{1}{2}qS^{-1}[r-St]\mod\frac{qN}{d}.$
The first property can be seen easily and the second follows from the fact that we already have $\displaystyle{D\equiv\frac{1}{2}qS^{-1}[r-S\lambda]\mod\frac{qN}{d}}$ and also we can check
So, it suffices to show that
$\frac{1}{2}S[t]\equiv\frac{1}{2}S[\lambda]\mod{\frac{N}{d}}.$
But this follows after we write $\lambda=t+N\alpha/d$ with $\alpha\in\mathbb{Z}^{n}$, from (6).
Hence, because of Remark 2, expression (5) becomes (after replacing $d$ with $N/d$)
$\sum_{d\mid N}d^{k-(n+1)}\sum_{t\in\mathbb{Z}^{n}/d\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r-St]\mod qd\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}(r-St)\right)\times\\ \times e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right).$
What is left to prove now (after fixing $D,r$ with the appropriate conditions) is
$\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}t\mod dS\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[t]\mod{qd}\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}t\right).$
This follows by setting $u=r-St$ and then observing that $\displaystyle{D\equiv\frac{1}{2}qS^{-1}[u]\mod qd}$,
and by using the fact that $c_{\phi}(D,s)=c_{\phi}(D,s^{\prime})$ if $s\equiv s^{\prime}\mod NS\mathbb{Z}^{n}$ (see Remark 2). So we can consider the entries $r-St$ $\ (\mathrm{mod}\ NS\mathbb{Z}^{n})$ and all of these are different $\ (\mathrm{mod}\ dS\mathbb{Z}^{n})$.∎
@@ 5. Poincaré Series
In this Section, we define a special class of Fourier–Jacobi forms, called Poincaré series, which reproduce the Fourier coefficients of Jacobi forms under the Petersson scalar product. What is more, they generate the space of cusp forms and in turn the Maass space, as we will show in the next Section.
@@
We define the support of the lattice $L=\mathbb{Z}^{n}$ with respect to the bilinear form $\sigma(x,y)=x^{t}Sy$ for $x,y\in V$ to be
@@
which then implies $D\in\frac{1}{q}\mathbb{Z}$. Hence, by writing $D\longmapsto D/q$, we get that equivalently
@@
$\widetilde{\textup{supp}}(L,\sigma):=\{(D,r)\in\mathbb{Z}_{\leq 0}\times L\mid(D/q,S^{-1}r)\in\textup{supp}(L,\sigma)\}.$
@@
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. We then define the following complex valued function on $\mathbb{H}\times(L\otimes\mathbb{C})$:
$g_{D,r}(\tau,z):=e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right),$
where $q$ is the level of $L$ (see Definition 2.4).
@@
If $k>n/2+2$, then $P_{k,D,r}$ is absolutely and uniformly convergent on compact subsets of $\mathbb{H}\times(L\otimes\mathbb{C})$ and defines an element of $S_{k}(L,\sigma)$ (see [17, Theorem 2.3, (i)]). Moreover, from the same Theorem, it only depends on $S^{-1}r\pmod{L}$.
We now have the following very important property, which can be again found in [17, Theorem 2.3, (i)].
@@
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. Then for any $f\in S_{k}(L,\sigma)$ with
@@
6. Dirichlet Series and relation with Fourier coefficients
In this Section, we define the Dirichlet series of interest. Let $\widetilde{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$ be a subgroup. Let $F,G\in S_{k}(\Gamma)$ with corresponding Fourier–Jacobi coefficients $\{\phi_{m}\}_{m=1}^{\infty},\{\psi_{m}\}_{m=1}^{\infty}$ respectively (see (3)). We define the Dirichlet series
$\mathcal{D}_{F,G}(s):=\sum_{N=1}^{\infty}\langle\phi_{N},\psi_{N}\rangle N^{-s},$ (7)
where $\langle\ ,\ \rangle$ denotes the inner product of Definition 3.5. We then have the following Lemma.
Lemma 6.1.
$\mathcal{D}_{F,G}(s)$ converges absolutely for $\textup{Re}(s)>k+1$ and represents a holomorphic function in this domain.